Contact Us Follow us on Twitter Follow us on Linkedin

0 Open      0 Scheduled      82 Resolved

Spamexperts Reboot (Resolved)

Affecting System - spamexperts
Date - 08/08/2017 19:30 - 08/08/2017 21:30 Priority - High
During the scheduled window, the spamexperts appliances will be rebooted for updates to be applied.  They will be rebooted one at a ttime until all finished and no impact should happen to mail delivery as they are setup in a failover setup.  If you have any questions regarding this, please let us know.

Work has been completed.
Last Updated - 08/08/2017 18:48

mail1 outage (Resolved)

Affecting Server - Hsphere-Linux
Date - 03/02/2017 15:48 Priority - Critical
Due to a network issue on mail1, services are currently disrupted.  The issue is being investiaged at this time.

4:51 PM ET - issue has been resolved.
Last Updated - 03/02/2017 15:52

mail7-net (Resolved)

Affecting Server - Hsphere-Windows
Date - 02/23/2017 16:51 Priority - Critical
We are aware of service disruption with mail7-net, used by windows only accounts.   Further updates will be posted once more info is available.

7:50 PM ET - server is under a file system check, which unfortunately takes a while as mail servers hae a lot of small files and all have to be checked.
10:30 PM ET - unfortunately the file system check is still going.
2:10 AM ET  - services are coming up at this point, however, final checks are continuing.
2:30 AMT ET  - there still some work on this server at this time.
3 AM ET - services have been functional for the last 20 minutes or so and this outage should be resolved at this point.
4 AM ET - unfortunately the server is experiencing addtional outages at this point. Techs are working on this as quick as possible.
4:30 AM ET - hardware faults are being diagnosed at this point
6:00 AM ET - hardware faults have been resolved, however, some system software faults are now coming up preventing a clean bootup process and that is being worked on currently
7:30 AM ET - server is back in operation at this point.  We are sorry for the extended outage.
Last Updated - 02/24/2017 06:50
Affecting Server - Hsphere-Linux
Date - 10/01/2016 13:00 - 10/01/2016 15:00 Priority - Critical

As previosuly announced, we will be removing PHP 5.2 from our shared linux servers and the default PHP version for the near future will be PHP 5.3 with that planned to be removed later this year or next as well. 

This should not affect most sites, however, if you are using a database and after the removal you receive a database connection error, please make sure to update the database user's password using the long hash option.   Please see https://u2-web.com/knowledgebase/146/Update-MySQL-User-Password-To-Use-Long-Hash.html for more information.


Update:  This has been pushed to 10/3/2016 in order to prevent issues over the weekend. 


10/3/2016 - 8 PM ET:  php 5.2 has been removed from web2-dal1 web server.  Auto checks for sites is in progrses at this time.  After that is done, web1-dal1 will be worked on.  To see which web server you are on, check the control panel -> FTP user / account -> web server. 
10/3/2016- 8:30 PM:  web2-dal1 Auto checks of sites returned a few sites with database connection errors and those were fixed.  The checks only checked for keywords on the base of the site and not the whole site.  Again, if you see database errors on your sites, blogs, etc please update the mysql's password hash. 

Moving on to web1-dal1 and will update here once done.

10/3/2016 9:35 PM ET - PHP 5.2 has been removed from all shared linux web servers at this point.  Default PHP version is now 5.3 for the near future.  Sites have had a rough auto check for usual database connection errors and those found have been fixed.  However, this does not include sub addresses such as /blog or scripts that do not provide an error.  Again, if you are seeing database connection errors, please update the mysql password hash as described in https://u2-web.com/knowledgebase/146/Update-MySQL-User-Password-To-Use-Long-Hash.html

If you have any questions or issues, please contact support -- https://u2-web.com/contact/
This notice will be left open for at least a few days.

Last Updated - 02/24/2017 05:45
Affecting Server - Hsphere-Linux
Date - 02/23/2017 19:00 - 02/23/2017 22:59 Priority - High
During the scheduled window, services for mail1 will be moved to newly deployed hardware node and running new version of the operating system.  This move is required to keep servcices updated and a in place upgrade is not possible.

Due to the volume of mail, mail files will be migrated eariy on the day of the window and then a resync afterwards.  Due to this, you may see mail that was deleted / moved during this period come back. 

Actual downtime is expected to be about 30 minutes or less, however, we have reserved an extended period in case of any issues. 

Further updates will be provided as work begins.

4:10 PM ET - Prep work, including the mentioned message file copying, is starting at this time.

7:40 PM ET - mail message file copying is still underway is taking longer than expected.  If the copying is not done by 10 PM ET, this work will be pushed to tomorrow night.

10:05 PM ET - message file copy is still going at this time .  It will likely not completel until late tonight and as such actual switching will not be done until tomorrow night.  No services are affected at this point.

7:25 PM ET 2/17/2017 - This work is being pushed until Saturday (tomorrow) due to other projects conflicting with the timing.

7:00 PM ET 2/18/2017 - The switch of servers is happening at this time.  Again, downtime is expected to be less than 30 minutes but could extend further if issues come up.
7:30 PM ET - unfortunately some issues have come up for this work and it is being worked on.  If not resolved soon, backpedaling will be done.
7:45 PM ET - due to issues in the final switch over, this work has been backpedaled to the previous server and services are up at this time.  We will schedule another attempt later this week.

2/22/2017 - Will be attempting the switch to the new server tomorrow night.  Issues during the last migration have been sorted and it should be a smooth switch though again extended window for the move is scheduled. 

2/23/2017 - 7:10 PM ET - this migration is starting.  Further updates will be done shortly.
7:40 PM ET - services have been up for about 20 minutes now without issue.  Tests show mail is being delivered properly.  Clean up work is under way at his point.
8:20 PM ET - mail1 is going down for a quick reboot.
8:26 PM ET _ mail1 is back in operation .
9:55 PM ET -  Finaly mail message sync is under way at this time.  After that, this work is completed.   If you notice any issues, please let us know via support tickets.
10:10 PM ET - mail sync is completed.  Old server has been shutdown.  This work is now completed.
Last Updated - 02/23/2017 21:08

mail2 migration (Resolved)

Affecting Server - Hsphere-Linux
Date - 10/19/2016 21:00 - 10/19/2016 23:59 Priority - High
During the scheduled window we will be migrating mail2 to a new operating system install in order to keep it updated .  A in place update to the new system is not possible due to the operating system version differences.   During this process mail will be sync'd from the old / current system to the new one.  Once the switch is done a furthe sync will be done to grab any mail that came in during the first sync.  As such, you may see mail that you deleted on 10/18/2016 come back once the switch has been completed. Unfortunately there is no around that.

There is no action needed on customers part.   Accounts can have a mix of the mail servers we have due to the clustering nature.

This process may require an hour downtime for a final sync.  If that is required, any incoming mail will be queued at the sending mail server and attempted again per the sending mail server's settings (which can vary). 

Update:  10/18/2016 - 11 AM ET - Preparations, including initial sync of mail files, has started for this work.  As it can take several hours for the transfer of all the mail files, it is best to start now so that the switch can happen as scheduled.  No downtime should happen during this time.

Update: 10/18/2016 11 PM ET -- Due to issues on final prep work before switch, this work is being postpoined until tomorrow night.

Update: 10/19/2016 8:30 PM ET - We are on track for the scheduled switch in 30 minutes.

Update: 10/19/2016 9 PM ET - mail2 is being taken down at this time for the switch to the new server.

Update: 10/19/2016 21:25 PM ET -- swithch as been completed.  Downtime was around 5 minutes .  A final sync of mail is in progress.

Update: 10/29/2016 21:40 PM ET - sync is done and this work is now complete.  Tests show mail is being delivered as normal and mail is being sent out as expected.
Last Updated - 10/19/2016 20:42
Affecting Server - Hsphere-Linux
Date - 10/12/2016 21:00 - 10/13/2016 00:00 Priority - Critical
During the scheduled window, updates will be applied to the shared linux servers.  A reboot of the server may be needed.  Downtime is expected to be about 10 to 20 minutes per server.

Updates have been applied.  A reboot was not eeded at this time.  There was no downtime.
Last Updated - 10/12/2016 20:48

Win1 Updates (Resolved)

Affecting Server - Hsphere-Linux
Date - 09/28/2016 21:00 - 09/28/2016 00:00 Priority - High
During the scheduled window, updates for Windows and other software will be installed.   A reboot or two is expected to be needed due to the updates. 
  • Update Windows packages
  • Update control panel packages
  • Update Anti-virus software

Downtime is expected to be 10 to 20 minutes.


9:00 PM ET:  Updates are being applied at this time.
9:37 PM ET:  server is rebooting now.
9:44 PM ET:  Server is booting up
9:46 PM ET:  Update configurations are being applied.  Web services are coming back online.
9:58 PM ET:  This work has been completed.  Downtime was about 10 minutes.
Last Updated - 09/28/2016 20:58
Affecting Server - Hsphere-Linux
Date - 09/20/2016 09:42 Priority - Critical
The hsphere control panel is currently offline due to a database issue with it.  It is currently running through a "garabage" collection process that requires the database to be offline.  This process usually runs automatically, however, it appears to have not done so for some time.   This process has been stalling all day and we have had to restart it multiple times.  The process is required as to prevent data loss.   Unfortunately, the process does not provide any progress report and so we can not provide an ETA on this.  

Account services are not affected, beyond not being able to manage their settings. 

We are looking at ways to either bypass this or speed it up. 

We are sorry for the inconvenience this may cause.

9/21/2016 - 02:00 -  we are able to bring the database online and the control panel is responding at this time.  However, an online vacuum has been started to clean up the database further so that this issue does not happen again.   As such, the performance of the control panel may be affected due to this process. 

Again, we apologize for the control panel issues.   Investigations on why this happened in the first place is underway in the background as well as other plans for the control panel in general.
Last Updated - 09/21/2016 00:53
Affecting Server - Hsphere-Linux
Date - 09/05/2016 20:00 - 09/05/2016 23:30 Priority - High
Updates will be applied during the scheduled window for PHP and any system updates for our shared web servers.  A couple minutes downtime is expected while updates apply at the final stages.
Last Updated - 09/14/2016 16:58